# Bayesian Estimation and Tuning-Free Rank Detection for Probability Mass Function Tensors

### v1

@@ III-B2 Posterior distribution $q^{\ast}_{n,r}(\bm{a}_{n,r})$ of the factor matrix columns $\bm{a}_{n,r}$
From (14), we select terms with a dependence on $\bm{a}_{n,r}$ and then apply (18). This yields
$\sum_{n=1}^{N}\sum_{r=1}^{R}\log q_{n,r}^{\ast}(\bm{a}_{n,r})$ $=\sum_{n=1}^{N}\sum_{r=1}^{R}\Bigg{(}\sum_{t=1}^{T}\mathbb{E}[z_{r,t}]\log a_{n,r,y_{n,t}}$ (25)
$+\sum_{i_{n}=1}^{I_{n}}(\alpha_{n,r}-1)\log a_{n,r,i_{n}}\Bigg{)}+\mathrm{const}.$
In order to simplify (25), we need to introduce a summation over the index $i_{n}$ in the first term. Since $y_{n,t}\in\{1,\dotsc,I_{n}\}$, (25) can be equivalently expressed as
$\sum_{n=1}^{N}\sum_{r=1}^{R}\log q_{n,r}^{\ast}(\bm{a}_{n,r})$ $=$ (26)
$\sum_{n=1}^{N}\sum_{r=1}^{R}\sum_{i_{n}=1}^{I_{n}}\Bigg{(}\Big{(}\sum_{t\in\Omega_{n,i_{n}}}\mathbb{E}[z_{r,t}]\Big{)}\log a_{n,r,i_{n}}$
$+(\alpha_{n,r}-1)\log a_{n,r,i_{n}}\Bigg{)}+\mathrm{const},$
where $\Omega_{n,i_{n}}\triangleq\{t:y_{n,t}=i_{n}\}\subset\{1,\dotsc,T\}$ is the set of indices $t$ in which the observation $y_{n,t}$ equals the discrete value $i_{n}$. Restricting the summation range of the index $t$ in this way ensures that (26) and (25) are equivalent.
We can now isolate the $(n,r)$-th distribution, factor out $\log a_{n,r,i_{n}}$, and take exponentials on both sides, giving
$q^{\ast}_{n,r}(\bm{a}_{n,r})\propto\prod_{i_{n}=1}^{I_{n}}a_{n,r,i_{n}}^{(\alpha_{n,r}+\sum_{t\in\Omega_{n,i_{n}}}\mathbb{E}[z_{r,t}])-1},~{}~{}\forall n,r.$ (27)
@@
$\widetilde{\alpha}_{n,r,i_{n}}\triangleq\alpha_{n,r}+\sum_{t\in\Omega_{n,i_{n}}}\mathbb{E}[z_{r,t}],~{}i_{n}=1,\dotsc,I_{n}.$ (28)
@@
$q^{\ast}_{n,r}(\bm{a}_{n,r})=\mathrm{Dir}(\bm{a}_{n,r}\,|\,\bm{\alpha}_{n,r})=C(\widetilde{\bm{\alpha}}_{n,r})\prod_{i_{n}=1}^{I_{n}}a_{n,r,i_{n}}^{\widetilde{\alpha}_{n,r,i_{n}}-1},$ (29)
@@
$\widehat{a}_{n,r,i_{n}}=\mathbb{E}_{q^{\ast}_{n,r}(\bm{a}_{n,r})}[a_{n,r,i_{n}}]=\frac{\widetilde{\alpha}_{n,r,i_{n}}}{\sum_{j=1}^{I_{n}}\widetilde{\alpha}_{n,r,j}},~{}\forall i_{n}.$ (30)
@@ III-B3 Posterior distribution $q^{\ast}_{z}(\bm{z}_{t})$ of the latent variable $\bm{z}_{t}$
$\log q^{\ast}_{z}(\bm{z}_{t})=\sum_{r=1}^{R}z_{r,t}\Big{(}\log\widetilde{\lambda}_{r}+\sum_{n=1}^{N}\log\widetilde{a}_{n,r,y_{n,t}}\Big{)}$ (31)
where $\log\widetilde{x}\triangleq\mathbb{E}[\log x]$ and the expectations are taken with respect to the variational distributions $q^{\ast}_{\lambda}(\bm{\lambda})$ and $q^{\ast}_{n,r}(\bm{a}_{n,r})$, respectively. Therefore, the expectations are given by (cf. (10))
$\log\widetilde{\lambda}_{r}$ $=\psi(\widetilde{\alpha}_{\lambda,r})-\psi\Big{(}\sum_{j=1}^{R}\widetilde{\alpha}_{\lambda,r}\Big{)},$ (32)
$\log\widetilde{a}_{n,r,y_{n,t}}$ $=\psi(\widetilde{\alpha}_{n,r,y_{n,t}})-\psi\Big{(}\sum_{j=1}^{I_{n}}\widetilde{\alpha}_{n,r,j}\Big{)}.$
@@
$\gamma_{r,t}\triangleq\exp\Big{\{}\log\widetilde{\lambda}_{r}+\sum_{n=1}^{N}\log\widetilde{a}_{n,r,y_{n,t}}\Big{\}}.$ (34)
Normalizing (33) gives us the optimal variational distribution as
@@
is the local variational parameter (for the local parameter $z_{r,t}$). Recall that $\bm{z}_{t}$ is one-hot encoded (see Section III-A). Thus, taking the expectation with respect to (35) reveals that
@@
and this expectation can be inserted into (22) and (28) to compute the respective Dirichlet parameters.
Algorithm 1. VB-PMF Input: Dataset $\bm{Y}$, initial rank $R$ Output: CPD estimates $\widehat{\bm{\lambda}}$ and $\{\widehat{\bm{A}}_{n}\}_{n=1}^{N}$, rank estimate $\widehat{R}$ 1: Set prior hyperparameters $\alpha_{\lambda}$ and $\alpha_{n,r},\,\forall n,r$. 2: Initialize $\widetilde{\bm{\alpha}}_{\lambda}$ and $\widetilde{\bm{\alpha}}_{n,r},\forall n,r$ randomly. 3: repeat 4:     Update $\rho_{r,t},\,\forall r,t$ using (32), (34), and (36). 5:     Update $\widetilde{\alpha}_{\lambda,r},\,\forall r$ using (22). 6:     Update $\widetilde{\alpha}_{n,r,i_{n}},\,\forall n,r,i_{n}$ using (28). 7:     Evaluate the ELBO using (39). 8: until the ELBO converges 9: Find $\widehat{\bm{\lambda}}$ and $\{\widehat{\bm{A}}_{n}\}_{n=1}^{N}$ using (24) and (30), respectively. 10: Find $\widehat{R}$ by deleting components corresponding to $\quad\widehat{\lambda}_{r}<\epsilon$.
@@ III-C Computing the ELBO
As mentioned earlier in Section III-B, the goal of variational inference is to find a distribution $q(\bm{\Theta})$ that is closest to the true posterior distribution $p(\bm{\Theta}\,|\,\bm{Y})$ in terms of the KLD. According to (16), instead of directly optimizing the KLD, we only need to maximize the ELBO $\mathcal{L}(q)$. Since $\mathcal{L}(q)$ is a lower bound, it should increase at each iteration during optimization and can therefore be used to evaluate the correctness of the mathematical expressions, as well as to test for convergence. From (16), the ELBO can be written as
@@
where the first term denotes the posterior expectation of the joint distribution (cf. (14)), while the second term is the (negative) entropy of the posterior distributions. The expectations are taken with respect to the optimal variational distributions in turn. For the first term in (38), the joint distribution $p(\bm{Y},\bm{\Theta})$ can be decomposed into various terms as shown in (13). Evaluating the expectations yields the closed-form expression in (39) for computing the ELBO. Details on the derivation of (39) can be found in the appendix.
$\mathcal{L}(q)$ $=\sum_{t=1}^{T}\sum_{r=1}^{R}\rho_{r,t}\Big{(}\log\widetilde{\lambda}_{r}+\sum_{n=1}^{N}\log\widetilde{a}_{n,r,y_{n,t}}\Big{)}+\sum_{n=1}^{N}\sum_{r=1}^{R}\Big{(}\log C(\bm{\alpha}_{n,r})+(\alpha_{n,r}-1)\sum_{i_{n}=1}^{I_{n}}\log\widetilde{a}_{n,r,i_{n}}\Big{)}$ (39)
$+\Big{(}\log C(\bm{\alpha}_{\lambda})+(\alpha_{\lambda}-1)\sum_{r=1}^{R}\log\widetilde{\lambda}_{r}\Big{)}-\sum_{t=1}^{T}\sum_{r=1}^{R}\rho_{r,t}\log\rho_{r,t}-\Big{(}\log C(\widetilde{\bm{\alpha}}_{\lambda})+\sum_{r=1}^{R}(\widetilde{\alpha}_{\lambda,r}-1)\log\widetilde{\lambda}_{r}\Big{)}$
$-\sum_{n=1}^{N}\sum_{r=1}^{R}\Big{(}\log C(\widetilde{\bm{\alpha}}_{n,r})+\sum_{i_{n}=1}^{I_{n}}(\widetilde{\alpha}_{n,r,i_{n}}-1)\log\widetilde{a}_{n,r,i_{n}}\Big{)}$
@@ III-D Algorithm: VB-PMF
where we have used (22) and (37) together with the fact that $\sum_{t,r}\rho_{r,t}=T$. If the $r$-th component is not responsible for explaining the observations, then $\sum_{t}\rho_{r,t}\approx 0$. Therefore, for such a component, if $\alpha_{\lambda}\rightarrow 0$, then $\widehat{\lambda}_{r}\approx 0$ and the entire rank-one component (i.e., $\widehat{\lambda}_{r}$ and $\{\widehat{\bm{a}}_{n,r}\}_{n=1}^{N}$) can be pruned out of the model. The rank can therefore be determined via a single training run in which the algorithm is initialized with $R>R_{\mathrm{true}}$ components (where $R_{\mathrm{true}}$ is the unknown true rank of the PMF tensor), followed by pruning out components with a value less than some specified small positive constant $\epsilon$ after convergence. One strategy for choosing the initial number of components is to set $R$ to be the maximum possible rank for which the Kruskal uniqueness conditions for the CPD are satisfied [26], [27]. Alternatively, the initial rank can be set manually for computational efficiency. The algorithm is deemed to have converged when the difference between successive ELBO values is less than a specified threshold. Alternatively, instead of computing the ELBO of the full dataset, which could be expensive, the log-likelihood of a small held-out dataset can be used to test for convergence.
The complexity of the VB-PMF algorithm consists of updating the parameters of the variational distributions and computing the ELBO to check for convergence. The parameters $\rho_{r,t}$ are computed in $\mathcal{O}(NRT)$ while the combined complexity of computing $\widetilde{\bm{\alpha}}_{\lambda}$ and $\widetilde{\bm{\alpha}}_{n,r}$ is $\mathcal{O}(NR)$. In addition, the complexity of computing the ELBO is $\mathcal{O}(NRT)$. Thus, the overall complexity is $\mathcal{O}(NRT)$.
IV Scalable Bayesian PMF Estimation
In Section III, we have presented a PMF estimation approach based on variational Bayesian inference. One drawback of the VB-PMF algorithm is that, due to the coordinate ascent structure, the entire dataset needs to be analyzed at each iteration before updating the global variational parameters $\widetilde{\bm{\alpha}}_{\lambda}$ and $\widetilde{\bm{\alpha}}_{n,r}$. Thus, as the size of the dataset grows, VB-PMF does not scale well because each iteration becomes more computationally expensive. Scalability is especially important in the PMF estimation problem when the number of variables is large since a lot of data is required to obtain a reasonably accurate estimate of the joint PMF (even more if there are missing observations in the data). In this section, we present an improvement to VB-PMF which allows for scalability to massive datasets. The proposed algorithm is based on stochastic variational inference (SVI) [22], a method that scales variational inference by employing stochastic gradient-based optimization techniques in order to maximize the ELBO.
IV-A Preliminaries
We begin by briefly reviewing the exponential family of distributions (see, e.g., [1]). Given a random vector $\bm{x}\in\mathbb{R}^{n}$ governed by parameters $\bm{\theta}\in\mathbb{R}^{n}$, the exponential family of distributions over $\bm{x}$ given $\bm{\theta}$ is of the form
$p(\bm{x}\,|\,\bm{\theta})=h(\bm{x})g(\bm{\theta})\exp\{\bm{\eta}\!^{\mathsf{T}}\!(\bm{\theta})\bm{u}(\bm{x})\},$ (41)
where $h(\bm{x}):\mathbb{R}^{n}\mapsto\mathbb{R}$ is the base measure, $\bm{u}(\bm{x}):\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ are the sufficient statistics, $\bm{\eta}(\bm{\theta}):\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ are the natural parameters, and $g(\bm{\theta}):\mathbb{R}^{n}\mapsto\mathbb{R}$ is the normalization constant which ensures that $\int p(\bm{x}\,|\,\bm{\theta})\mathrm{d}\bm{x}=1$. Many of the most common probability distributions (e.g., normal, exponential, Bernoulli, Dirichlet, among others) are members of the exponential family. One attractive property of the exponential family is that certain moments can be easily calculated using differentiation. For instance, it can be shown that the expected value and the covariance of the sufficient statistics are given by
$\mathbb{E}[\bm{u}(\bm{x})]=-\nabla\log g(\bm{\theta}),$ (42)
$\mathrm{cov}[\bm{u}(\bm{x}),\bm{u}(\bm{x})]=\nabla^{2}\log g(\bm{\theta}).$ (43)
In addition, since $\bm{u}(\bm{x})$ are sufficient statistics, then (43) is also the Fisher information provided by the sample $\bm{x}$ about the parameter $\bm{\theta}$.
We are interested in using SVI to optimize the global variational parameters $\widetilde{\bm{\alpha}}_{\lambda}$ and $\widetilde{\bm{\alpha}}_{n,r}$, whose optimal variational distributions were found to be Dirichlet distributions (see (23) and (29)). It is useful to express the Dirichlet distribution in exponential form as such representation will come in handy in the sequel. Given a general Dirichlet distribution over a random vector $\bm{\mu}\in\mathbb{R}^{K}$ with parameters $\bm{\alpha}\in\mathbb{R}^{K}$, the exponential form is written as (cf. (7))
$\mathrm{Dir}(\bm{\mu}\,|\,\bm{\alpha})=h(\bm{\mu})g(\bm{\alpha})\exp\{\bm{\eta}\!^{\mathsf{T}}\!(\bm{\alpha})\bm{u}(\bm{\mu})\},$ (44)
with $h(\bm{\mu})=1$, $g(\bm{\alpha})=C(\bm{\alpha})$, $\bm{\eta}(\bm{\alpha})=[\alpha_{1}-1,\dotsc,\alpha_{K}-1]^{\mathsf{T}}$, and $\bm{u}(\bm{\mu})=[\log\mu_{1},\cdots,\log\mu_{K}]^{\mathsf{T}}$.
In the next subsection, we will use the exponential forms and the properties thereof to derive expressions for the natural gradient of the ELBO. These expressions will subsquently play an important role in the development of an SVI-based algorithm.
IV-B Computing the natural gradient of the ELBO
In classical gradient-based optimization where the goal is to maximize an objective function, the algorithm takes steps in the direction that results in the largest value for the function, but also maintains the shortest Euclidean distance between the steps (i.e., the direction of steepest ascent). However, the Euclidean distance does not capture the notion of (dis)similarity between probability distributions meaningfully; a more appropriate metric is the symmetric KLD, defined as $D(p\,\|\,q)+D(q\,\|\,p)$ for two distributions $p$ and $q$. Let the optimization problem be defined such that the steps taken minimize the symmetric KLD (as opposed to the Euclidean distance). The gradient which points in the direction of steepest ascent in the symmetric KLD metric space is termed the natural gradient [28].
Consider the parameter vector $\bm{\theta}$ defined in (41) and assume its assigned prior distribution is in the exponential family, i.e., $p(\bm{\theta}\,|\,\bm{\beta})=h(\bm{\theta})g(\bm{\beta})\exp\{\bm{\eta}\!^{\mathsf{T}}\!(\bm{\beta})\bm{u}(\bm{\theta})\}$. Suppose that we have approximated the true posterior distribution $p(\bm{\theta}\,|\,\bm{x})$ using a variational distribution $q^{\ast}_{\theta}(\bm{\theta}\,|\,\widetilde{\bm{\beta}})$. Our goal is to maximize the ELBO $\mathcal{L}(\widetilde{\bm{\beta}})$ over the variational parameter $\widetilde{\bm{\beta}}$. It has been shown in [22] that if $q^{\ast}_{\theta}(\bm{\theta}\,|\,\widetilde{\bm{\beta}})$ is in the exponential family, then the natural gradient of the ELBO is given by
~$\widetilde{\nabla}\mathcal{L}(\widetilde{\bm{\beta}})=\big{(}-\nabla^{2}\log g(\widetilde{\bm{\beta}})\big{)}^{-1}\cdot\nabla\mathcal{L}(\widetilde{\bm{\beta}}),$ (45)
where $\nabla^{2}\log g(\widetilde{\bm{\beta}})$ is the Fisher information and $\nabla\mathcal{L}(\widetilde{\bm{\beta}})$ is the gradient of $\mathcal{L}(\widetilde{\bm{\beta}})$.
We are now ready to derive the natural gradients with respect to the global variational parameters $\widetilde{\bm{\alpha}}_{\lambda}$ and $\widetilde{\bm{\alpha}}_{n,r}$. The ELBO expressed as a function of $\widetilde{\alpha}_{\lambda,r}$ is given by (cf. (39))
$\mathcal{L}(\widetilde{\alpha}_{\lambda,r})=\mathbb{E}[\log\lambda_{r}]\Bigg{(}\alpha_{\lambda}+\sum_{t=1}^{T}\rho_{r,t}-\widetilde{\alpha}_{\lambda,r}\Bigg{)}+\mathrm{const},$ (46)
where we have explicitly written out the expected value in place of the notation $\log\widetilde{\lambda}_{r}$. Applying the exponential family property in (42) together with the general form in (44), we note that $\mathbb{E}[\log\lambda_{r}]=-\nabla_{{\widetilde{\alpha}_{\lambda,r}}}\log C(\widetilde{\bm{\alpha}}_{\lambda})$. Taking the gradient, we have
$\nabla_{\widetilde{\alpha}_{\lambda,r}}\mathcal{L}{\color[rgb]{0,0,0}(\widetilde{\alpha}_{\lambda,r})}=-\nabla_{\widetilde{\alpha}_{\lambda,r}}^{2}\log C(\widetilde{\bm{\alpha}}_{\lambda})\Bigg{(}\alpha_{\lambda}+\sum_{t=1}^{T}\rho_{r,t}-\widetilde{\alpha}_{\lambda,r}\Bigg{)},$ (47)
which is the gradient of the ELBO w.r.t the parameter $\widetilde{\alpha}_{\lambda,r}$. Finally, applying (45) with $\widetilde{\bm{\beta}}=\widetilde{\bm{\alpha}}_{\lambda}$ yields the natural gradient
~$\widetilde{\nabla}_{\widetilde{\alpha}_{\lambda,r}}\mathcal{L}{\color[rgb]{0,0,0}(\widetilde{\alpha}_{\lambda,r})}=\alpha_{\lambda}+\sum_{t=1}^{T}\rho_{r,t}-\widetilde{\alpha}_{\lambda,r},~{}\forall r.$ (48)
By following a similar procedure, the natural gradient with respect to the parameter $\widetilde{\alpha}_{n,r,i_{n}}$ can also be found as
~$\widetilde{\nabla}_{\widetilde{\alpha}_{n,r,i_{n}}}\mathcal{L}{\color[rgb]{0,0,0}(\widetilde{\alpha}_{n,r,i_{n}})}=\alpha_{n,r}+\sum_{t\in\Omega_{n,i_{n}}}\rho_{r,t}-\widetilde{\alpha}_{n,r,i_{n}},~{}\forall n,r,i_{n}.$ (49)
Notice the relationship between the natural gradients derived here and the variational update steps in (22) and (28), respectively. The natural gradient is therefore the difference between the variational updates and the current setting of the parameters.
Algorithm 2. SVB-PMF
Input: Dataset $\bm{Y}$, initial rank $R$, batch size $M$

### v2

@@ III-B2 Posterior distribution $q^{\ast}_{n,r}(\bm{a}_{n,r})$ of the factor matrix columns $\bm{a}_{n,r}$
From (14), we select the terms with a dependence on $\bm{a}_{n,r}$ and then apply (18). Isolating the $(n,r)$-th term, we have
$\log q_{n,r}^{\ast}(\bm{a}_{n,r})$ $=\sum_{t=1}^{T}\mathbb{E}[z_{r,t}]\log a_{n,r,y_{n,t}}$ (25)
$\hskip-20.00003pt+\sum_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}(\alpha_{n,r}-1)\log a_{n,r,{\color[rgb]{0,0,0}i}}+\mathrm{const}.$
To simplify (25), we rewrite the first term using a summation over the index $i$. Since $y_{n,t}\in\{1,\dotsc,I_{n}\}$, (25) can be equivalently expressed as
$\log q_{n,r}^{\ast}(\bm{a}_{n,r})$ $=\sum_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}\Bigg(\Big(\sum_{t\in\Omega_{n,{\color[rgb]{0,0,0}i}}}\mathbb{E}[z_{r,t}]\Big)\log a_{n,r,{\color[rgb]{0,0,0}i}}$ (26)
$+(\alpha_{n,r}-1)\log a_{n,r,{\color[rgb]{0,0,0}i}}\Bigg)+\mathrm{const},$
where $\Omega_{n,{\color[rgb]{0,0,0}i}}\triangleq\{t:y_{n,t}={\color[rgb]{0,0,0}i}\}\subset\{1,\dotsc,T\}$ is the set of indices $t$ in which the observation $y_{n,t}$ equals the discrete value $\color[rgb]{0,0,0}i$}. Restricting the summation range of the index $t$ in this way ensures that (26) and (25) are equivalent. The optimal variational distribution $q^{\ast}_{n,r}(\bm{a}_{n,r})$ is therefore given by
$q^{\ast}_{n,r}(\bm{a}_{n,r})\propto\prod_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}a_{n,r,{\color[rgb]{0,0,0}i}}^{(\alpha_{n,r}+\sum_{t\in\Omega_{n,{\color[rgb]{0,0,0}i}}}\mathbb{E}[z_{r,t}])-1},~~\forall n,r.$ (27)
@@
$\widetilde{\alpha}_{n,r,{\color[rgb]{0,0,0}i}}\triangleq\alpha_{n,r}+\sum_{t\in\Omega_{n,{\color[rgb]{0,0,0}i}}}\mathbb{E}[z_{r,t}],~{\color[rgb]{0,0,0}i}=1,\dotsc,I_{n}.$ (28)
@@
$q^{\ast}_{n,r}(\bm{a}_{n,r})=\mathrm{Dir}(\bm{a}_{n,r}\,|\,\bm{\alpha}_{n,r})=C(\widetilde{\bm{\alpha}}_{n,r})\prod_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}a_{n,r,{\color[rgb]{0,0,0}i}}^{\widetilde{\alpha}_{n,r,{\color[rgb]{0,0,0}i}}-1},$ (29)
@@
$\widehat{a}_{n,r,{\color[rgb]{0,0,0}i}}=\mathbb{E}_{q^{\ast}_{n,r}(\bm{a}_{n,r})}[a_{n,r,{\color[rgb]{0,0,0}i}}]=\frac{\widetilde{\alpha}_{n,r,{\color[rgb]{0,0,0}i}}}{\sum_{j=1}^{I_{n}}\widetilde{\alpha}_{n,r,j}},~\forall{\color[rgb]{0,0,0}i}.$ (30)
@@ III-B3 Posterior distribution $q^{\ast}_{z}(\bm{z}_{t})$ of the latent variable $\bm{z}_{t}$
$\log q^{\ast}_{z}(\bm{z}_{t})=\sum_{r=1}^{R}z_{r,t}\Big({\color[rgb]{0,0,0}\mathbb{E}[\log\lambda_{r}]}+\sum_{n=1}^{N}{\color[rgb]{0,0,0}\mathbb{E}[\log a_{n,r,y_{n,t}}]}\Big)$ (31)
where the expectations are taken with respect to the variational distributions $q^{\ast}_{\lambda}(\bm{\lambda})$ and $q^{\ast}_{n,r}(\bm{a}_{n,r})$, respectively. Therefore, the expectations are given by (cf. (10))
${\color[rgb]{0,0,0}\mathbb{E}[\log\lambda_{r}]}$ $=\psi(\widetilde{\alpha}_{\lambda,r})-\psi\Big(\sum_{j=1}^{R}\widetilde{\alpha}_{\lambda,r}\Big),$ (32)
${\color[rgb]{0,0,0}\mathbb{E}[\log a_{n,r,y_{n,t}}]}$ $=\psi(\widetilde{\alpha}_{n,r,y_{n,t}})-\psi\Big(\sum_{j=1}^{I_{n}}\widetilde{\alpha}_{n,r,j}\Big).$
@@
$\gamma_{r,t}\triangleq\exp\Big\{{\color[rgb]{0,0,0}\mathbb{E}[\log\lambda_{r}]}+\sum_{n=1}^{N}{\color[rgb]{0,0,0}\mathbb{E}[\log a_{n,r,y_{n,t}}]}\Big\}.$ (34)
Normalizing (33) gives us the optimal variational distribution
@@
is the local variational parameter (for the local parameter $z_{r,t}$). Recall that $\bm{z}_{t}$ is one-hot-encoded (see Section III-A). Thus, taking the expectation with respect to (35) reveals that
@@
and this expectation can be substituted into (22) and (28) to compute the respective Dirichlet parameters.
Table I summarizes the prior and posterior distributions, update equations, and point estimates for each model parameter. Note that the updates are easily adapted to datasets with missing observations (i.e., $y_{n,t}=0$; cf. (3)) by only considering sample indices $t\in\Omega_{n,i}$, i.e., those indices for which $y_{n,t}=i$, where $i=1,\dotsc,I_{n}$.
@@ III-C Computing the ELBO
As mentioned in Section III-B, the goal of variational inference is to find a distribution $q(\bm{\Theta})$ that is closest to the true posterior distribution $p(\bm{\Theta}\,|\,\bm{Y})$ in terms of the KLD. According to (16), rather than directly optimizing the KLD, we only need to maximize the ELBO $\mathcal{L}(q)$. Since $\mathcal{L}(q)$ is a lower bound, it should increase at each iteration during optimization and can therefore be used to evaluate the correctness of the mathematical expressions and to test for convergence. From (16), the ELBO can be written as
@@
where the first term denotes the posterior expectation of the joint distribution (cf. (14)), while the second term is the (negative) entropy of the posterior distributions. The expectations are taken with respect to the optimal variational distributions in turn. For the first term in (38), the joint distribution $p(\bm{Y},\bm{\Theta})$ can be decomposed into various terms, as shown in (13). Evaluating the expectations yields the closed-form expression in (39) for computing the ELBO. Details on the derivation of (39) can be found in the appendix.
@@ III-D Algorithm: VB-PMF
$\mathcal{L}(q)$ $=\sum_{t=1}^{T}\sum_{r=1}^{R}\rho_{r,t}\Big({\color[rgb]{0,0,0}\mathbb{E}[\log\lambda_{r}]}+\sum_{n=1}^{N}{\color[rgb]{0,0,0}\mathbb{E}[\log a_{n,r,y_{n,t}}]}\Big)+\sum_{n=1}^{N}\sum_{r=1}^{R}\Big(\log C(\bm{\alpha}_{n,r})+(\alpha_{n,r}-1)\sum_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}{\color[rgb]{0,0,0}\mathbb{E}[\log a_{n,r,{\color[rgb]{0,0,0}i}}]}\Big)$ (39)
$+\Big(\log C(\bm{\alpha}_{\lambda})+(\alpha_{\lambda}-1)\sum_{r=1}^{R}{\color[rgb]{0,0,0}\mathbb{E}[\log\lambda_{r}]}\Big)-\sum_{t=1}^{T}\sum_{r=1}^{R}\rho_{r,t}\log\rho_{r,t}-\Big(\log C(\widetilde{\bm{\alpha}}_{\lambda})+\sum_{r=1}^{R}(\widetilde{\alpha}_{\lambda,r}-1){\color[rgb]{0,0,0}\mathbb{E}[\log\lambda_{r}]}\Big)$
$-\sum_{n=1}^{N}\sum_{r=1}^{R}\Big(\log C(\widetilde{\bm{\alpha}}_{n,r})+\sum_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}(\widetilde{\alpha}_{n,r,{\color[rgb]{0,0,0}i}}-1){\color[rgb]{0,0,0}\mathbb{E}[\log a_{n,r,{\color[rgb]{0,0,0}i}}]}\Big)$
Algorithm: VB-PMF
Input: Dataset $\bm{Y}$, initial rank $R$